A firm’s production can be described with the following production function q = 6L2 -L3 + 20L, the amount of capital is fixed at 3 units and the total revenues are given by TR = 5q. Labour supply that the firm faces is given by the function w = 100.

• Decide whether the company is a perfect or imperfect competitor in the output market.

• Decide whether the company is a perfect or imperfect competitor in the labour market.

• Express the function MRPL , ARPL , MFCL , and AFCL .

• Determine how employees the firm is willing to employ. Next, determine the wage rate that the firm will pay for each unit of labour.


1
Expert's answer
2021-12-02T10:38:38-0500

Solution:

Production function (q) = 6L2 – L3 + 20L

MPL = "\frac{\partial Q} {\partial L}" = 12L – 3L2 + 20

The company is an imperfect competitor in the output market since it is a monopoly.

 

The company is an imperfect competitor in the labor market.

 

MRPL = MR x MPL

MR = "\frac{\partial TR} {\partial Q} = 5"

MRPL = 5 "\times" (12L – 3L2 + 20) = 60L – 15L2 + 100

 

ARPL = AR "\times" APL

AR = "\frac{TR}{Q} = \frac{5Q}{Q} = 5"


APL = "\frac{Q}{L}" = 6L2 – L3 + "\frac{20L}{L}" = 6L – L2 + 20

ARPL = 5 x (6L – L2 + 20) = 30L – 5L2 + 100

 

MFCL = MFC x L

MFC = "\frac{\partial TFC} {\partial Q}" = 12L – 3L2 + 20

MFCL = 100 "\times" (12L – 3L2 + 20) = 1200L – 300L2 + 2000

 

AFCL = AFC "\times" L

AFC = "\frac{TFC}{Q}" = 6L2 – L3 + "\frac{20L}{L}" = 6L – L2 + 20

AFCL = 100. (6L – L2 + 20) = 600L – 100L2 + 2000

 

Number of employees to employ = "\frac{TFC}{Q}"

TFC = 6(32) – 33 + 20(3) = 87

= "\frac{87}{3}" = 29

Number of employees to employ = 29 employees

Wage rate = "\frac{29}{3}" = 9.6